### v2

# Fabian Hinder, Valerie Vaquet, and Barbara Hammer

Abstract
Concept drift refers to the change of data distributions over time. While drift poses a challenge for learning models, requiring their continual adaption, it is also relevant in system monitoring to detect malfunctions, system failures, and unexpected behavior. In the latter case, the robust and reliable detection of drifts is imperative. This work studies the shortcomings of commonly used drift detection schemes. We show that they are prone to adversarial attacks, i.e., streams with undetected drift. In particular, we give necessary and sufficient conditions for their existence, provide methods for their construction, and demonstrate this behavior in experiments.
1 Introduction
Data from the real world is often subject to continuous changes known as concept drift [1, 2, 3]. Such can be caused by seasonal changes, changed demands, aging of sensors, etc. Concept drift not only poses a problem for maintaining high performance in learning models [2, 3] but also plays a crucial role in system monitoring [1]. In the latter case, the detection of concept drift is crucial as it enables the detection of anomalous behavior. Examples include machine malfunctions or failures, network security, environmental changes, and critical infrastructures. This is done by detecting irregular drifts [4, 1, 5]. In these contexts, the ability to robustly detect drift is essential.
In addition to problems such as noise and sampling error, which challenge all statistical methods, drift detection faces a special kind of difficulty when the drift follows certain patterns that evade detection. In this work, we study those specific drifts that we will refer to as “drift adversarials”. Similar to adversarial attacks in classification, that exploit model properties to force wrong decisions [6], drift adversarials exploit weaknesses in the detection methods, and thus allow significant concept drift to occur without triggering alarms posing major issues for monitoring systems. Besides the construction of drift adversarials, the presented theory also provides tools to check whether a specific drift detector is provably correct.
This paper is structured as follows: First (Section 2) we recall the definition of concept drift and define the two setups for which we will construct our adversarials. In Section 3, we construct drift adversarials that exploit which data is used for the analysis. Here, we mainly focus on two window approaches. In the last part (Section 4) we perform a numerical evaluation of our considerations and conclude the work (Section 5).
Algorithm 1 Two Window Drift Detector (no memory management)
1:procedure DriftDetection($(x_{i})_{i=1}^{n}$ data stream, $\theta$ detection threshold)
2:    for $(W_{1},W_{2})\in\mathcal{W}(n)$ $\triangleright$ iterate over all considered window pairs
3:         $p\leftarrow\textsc{Test}(\{x_{i}\mid i\in W_{1}\},\{x_{i}\mid i\in W_{2}\})$triangleright Test samples in $W_{1}$ and $W_{2}$
4:         if $p<\theta$ then
5:             Alert drift
6:         end if
7:    end for
8:end procedure
2 Concept Drift and Drift Detection
Most machine learning research focuses on the batch setup where one considers a fixed data set as i.i.d. random variables $X_{1},\dots,X_{n}$ following some distribution $\mathcal{D}$ on the data space $\mathcal{X}$. However, in many scenarios, data is obtained as a stream over time and is thus prone to potential changes of the underlying distribution, a phenomenon known as concept drift [1, 3]. In such finite sample setup, drift is typically defined in a sample-wise sense, that is two samples not having the same distribution, i.e., $\exists i,j:\mathbb{P}_{X_{i}}\neq\mathbb{P}_{X_{j}}$ [3]. Drift detection refers to the task of deciding whether or not the stream is affected by drift. One issue of this setup is that one cannot estimate the distribution $\mathbb{P}_{X_{i}}$ the single sample $X_{i}$ follows [1].
To analyze this task theoretically, we consider an extension building on distribution processes [1] describing the limiting case. We model a time $\mathcal{T}$ indexed family of probability measures $\mathcal{D}_{t}$ on $\mathcal{X}$ together with an observation probability $P_{T}$ on $\mathcal{T}$ [1]. A stream consists of dated data points $(X_{1},T_{1}),(X_{2},T_{2}),\dots$ such that a data point $X_{i}$ observed at time $t$ follows the distribution $\mathcal{D}_{t}$, i.e., $T_{i}\sim P_{T}$ and $X_{i}\mid T_{i}=t\sim\mathcal{D}_{t}$. Concept drift occurs if the chance of observing two different distributions is larger zero [1], i.e., $\mathbb{P}[\exists i,j:\mathbb{P}_{X_{i}}\neq\mathbb{P}_{X_{j}}]>0$. Notice that this definition is not limited to abrupt drift but all kinds of drift, such as gradual or recurring, and the statistical nature resolves the estimation problem.
In this paper, we are interested in constructing scenarios containing drift that is not detected. We will first investigate this task theoretically by examining the limiting case leveraging the definition by distribution processes. This also allows us to prove guarantees. Afterward, we study the finite case and derive a practical algorithm for the construction of drift adversarials.
3 Adversarial Attacks for Drift Detection
Most drift detectors process data on sliding windows using some statistical tool, most commonly a metric [1] (see Algorithm 1). This allows for two natural attack scenarios: Metric Adversarials construct distributions indistinguishable by the metric, while Window Adversarials exploit the data selection stage.
3.1 Metric Adversarials
The most commonly used drift detectors are based on learning models referring to the optimal model or model accuracy to detect drift [3]. However, as already pointed out in [7] this approach is flawed and can be exploited in many cases. Indeed, the authors provide a constructive proof that can easily be modified to construct a metric adversarial. Other approaches for which metric adversarials can be constructed include methods like the windowing Kolmogorov-Smirnov test that operates feature-wise and thus ignores drifts in correlations as shown in [1] or methods that use deep embeddings for which classical adversarials can be constructed. However, in many cases, it is not possible to construct a metric adversarial, e.g., when the used metric is indeed a metric. We will therefore mainly focus on window adversarials in this paper.
3.2 Window Adversarials for Two-Window-Based Detectors
As most drift detectors work by comparing data from two windows [1] we will focus on this setup. In the following, we consider the limiting and finite case.
The limiting case
We refer to the case where we take the sampling rate to infinity so that errors due to sampling vanish and the drift detector becomes a map of the kernel $\mathcal{D}_{t}$, i.e., the limiting case of 3 takes on the form
$A(\mathcal{D}_{t})=\mathbf{1}\left[\sup_{(W_{1},W_{2})\in\mathcal{W}}d(\mathcal{D}_{W_{1}},\mathcal{D}_{W_{2}})>0\right]$ (1)
Table 1: Overview of improper adversarial functions for common windowing schemes used in drift detection (assuming Lebesgue measure $P_{T}=\lambda$). Cases with Boundary Effects (BE) are marked. Proofs in appendix.
$\mathcal{T}$ $\mathcal{W}$ $\textnormal{Adv}_{0}$ (Theorem 1) BE
$\mathbb{R}$ $([t-l,t],[t,t+l])$ $t\in\mathcal{T}$ $f(t)=f(t+l)$
two sliding windows $l$-periodic functions
[…]
fixed reference window $l$-periodic after $a$ with same mean
$\mathbb{R}_{\geq 0}$ $([0,t],[t,t+l])$ $t\geq a$ $a^{-1}\int_{0}^{a}f(s)\textnormal{d}s=f(a)=f(t)\forall t\geq a$
growing reference window arbitrary before $a$ and then constant
where $\mathcal{W}$ is the set of all window pairs directly compared by the detector and $d$ is the used metric. Here, $A$ detects drift if $A(\mathcal{D}_{t})=1$. As $A$ cannot have false positives, the window adversarials are given by false negatives which can be constructed as follows:
Theorem 1.
Define the improper adversarial functions for $A$ as in Eq. 1 as $\textnormal{Adv}_{0}(A)=\left\{f:\mathcal{T}\to[0,1]\>\left|\>P_{T}(W_{2})\int_{W_{1}}\!\!\!\!f\,\,\textnormal{d}P_{T}=P_{T}(W_{1})\int_{W_{2}}\!\!\!\!f\,\,\textnormal{d}P_{T}\forall(W_{1},W_{2})\in\mathcal{W}\right.\right\}$ (2) then $A$ detects no drift, i.e., $A(\mathcal{D}_{t})=0$, if and only if $t\mapsto\mathcal{D}_{t}(S)\in\textnormal{Adv}_{0}(A)$ for all measurable $S\subset\mathcal{X}$.
Define the adversarial functions $\textnormal{Adv}(A)\subset\textnormal{Adv}_{0}(A)$ as those that are not constant. Then, $\textnormal{Adv}(A)$ describes all distribution processes with drift that is not detected by $A$. In particular, for $f\in\textnormal{Adv}(A)$ and distributions $P\neq Q$ on $\mathcal{X}$, $\mathcal{D}_{t}=f(t)P+(1-f(t))Q$ is a window adversarial, i.e., $\mathcal{D}_{t}$ has drift and $A(\mathcal{D}_{t})=0$. Conversely, for every window adversarial $A(\mathcal{D}_{t})=0$ we have $t\mapsto\mathcal{D}_{t}(S)\in\textnormal{Adv}_{0}(A)$. Therefore, if $\textnormal{Adv}(A)=\emptyset$ then $A$ detects every drift assuming $d$ is a metric.
The proof is given in the appendix.
We want to stress that the adversaials do not depend on the metric $d$ but only the considered windows and that the drift detector detects every possible drift if and only if $\textnormal{Adv}(A)=\emptyset$. This can for example be achieved by combining multiple drift detectors as $\textnormal{Adv}(\,(A,B)\,)\subset\textnormal{Adv}(A)\cap\textnormal{Adv}(B)$.
Most detectors use a sliding window for the current distribution of fixed length. There are three main strategies for the reference window: 1) fixed, 2) growing, and 3) sliding with fixed length [1]. Furthermore, there are two update strategies: Either the update is performed after every single data point, which in the limit is for every time point, or by considering chunks of data points. For the latter, we can hide arbitrary drifts within a chunk allowing for trivial adversarials. For point-wise updates and any of the aforementioned reference windows, we present the adversarial functions in Table 1.
The finite case
Analog to the limiting case we can also consider the case of finite samples $X_{1},\dots,X_{n}$. In this case, the windows refer to which samples are considered, i.e., $W_{1},W_{2}\subset[n]$. We will denote the set of all window pairs for $n$ samples by $\mathcal{W}(n)$ together with a normalized distance measure, e.g., a statistic test, and a decision threshold $\theta$ this leads to Algorithm 1. Usually, there is some memory management so that we do not have to store the entire stream which however depends on the windowing scheme $\mathcal{W}(n)$.
Algorithm 2 Construction of Drift Adversarials
1:function ConstructDriftAdversarial($P,Q$ sampling distributions, $\mathcal{W}(n)$ windowing scheme to be attacked)
2:    $\mathbf{W}_{n}\leftarrow[\textbf{1}]$
3:    for $(W_{1},W_{2})\in\mathcal{W}(n)$ do
4:         $\mathbf{W}_{n}\leftarrow\mathbf{W}_{n}+\left[|W_{1}|^{-1}\sum_{i=1}^{n}\mathbf{1}[i\in W_{1}]e_{i}-|W_{2}|^{-1}\sum_{i=1}^{n}\mathbf{1}[i\in W_{2}]e_{i}\right]$
5:    end for
6:    $v\leftarrow\textsc{Solve}(\mathbf{W}_{n}x=0)$triangleright Interpret $\mathbf{W}_{n}$ as a matrix
7:    $v\leftarrow\frac{v-\min_{i}v_{i}}{\max_{i}v_{i}-\min_{i}v_{i}}$
8:    $x\leftarrow[]$
9:    for $i=1,\dots,n$ do
10:         $x\leftarrow x+[\textsc{Sample}(v_{i}P+(1-v_{i})Q)]$
11:    end for
12:    return $x$
13:end function
We can encode the window selection $\mathcal{W}(n)$ into a single weight matrix $\mathbf{W}_{n}$ that encodes the pair $(W_{1},W_{2})$ as the vector $w=|W_{1}|^{-1}\sum_{i\in W_{1}}e_{i}-$$|W_{2}|^{-1}\sum_{i\in W_{2}}e_{i}$ where $e_{i}$ is the $i$-th coordinate vector. This representation is quite useful as it for example allows us to write the biased MMD – a kernel-based probability metric commonly used in drift detection [1] – of the $i$-th window as $(\mathbf{W}_{n}^{\top}K\mathbf{W}_{n})_{ii}$ where $K_{ij}=k(X_{i},X_{j})$ is the kernel matrix. For our purpose, it is useful as the kernel of $\mathbf{W}_{n}$ can be used to construct drift adversarials. To do so choose $v\in[0,1]^{n}$ with $\mathbf{W}_{n}v=0$ and then sample $X_{i}\sim v_{i}P+(1-v_{i})Q$. This idea is represented in Algorithm 2. If $v$ is not constant and $P\neq Q$ then the distributions differ for some $X_{i}$, i.e., there is drift in the sample-wise sense, the mean distributions of the samples in $W_{1}$ and $W_{2}$ however coincide for all $(W_{1},W_{2})\in\mathcal{W}(n)$ which is what Algorithm 1 line 3 is testing for. There are ways to increase the quality by choosing $v\in\{0,1\}^{n}$ or trying to avoid fast oscillations as such streams are similar to non-drifting streams. Notice, that there is a tight connection between the limiting and the finite setup which is given by sampling adversarial functions (Theorem 1) equidistant to obtain $v$. Yet, $\{v\mid\mathbf{W}_{n}v=0\}$ can be much larger than $\textnormal{Adv}(A)$ due to boundary effects (BE in Table 1).
Instead of comparing the mean distribution of two windows, some drift detectors – dubbed block-based in [1] – check for any kind of drift within a single window. Using similar techniques it can be shown that such detectors are not prone to window adversarial attacks. In the next section, we will test our theoretical observations empirically.
4 Empirical Evaluation
To evaluate our methodology we consider two empirical setups: a numerical analysis on synthetic data, and a showcase on data from critical infrastructure. [^2] The code can be found at https://github.com/FabianHinder/Drift-Adversarials
Table 2: Result of numerical analysis. $90\%/10\%$-quintile of obtained $p$-values (500 runs). Correct result is $p=0$, lining marks adversarials according to theory. The number in brackets is the length of the initial reference window.
Dataset / $\mathcal{W}$ fixed (100) fixed (150) grow (100) grow (150) slideing
Periodic $\underline{0.63/0.28}$ 0.00/$0.00$/0.000.00 / 0.00 0.00/$0.00$/0.000.00 / 0.00 0.00/$00/0.0$ 0.00 $\underline{0.39/0.27}$
Rand.Const (100) $\underline{0.59/0.31}$ $\underline{0.54/0.31}$ $\underline{0.39/0.27}$ $\underline{0.46/0.28}$ 0.00/$00/0.0$ 0.00
Rand.Const (150) 0.00/$00/0.0$ 0.00 $\underline{0.50/0.25}$ 0.00/$00/0.0$ 0.00 $\underline{0.45/0.24}$ 0.00/$00/0.0$ 0.00
Rand.Per. (100) $\underline{0.49/0.26}$ 0.00/$0.00$/0.000.00 / 0.00 0.00/$0.00$/0.000.00 / 0.00 0.00/$0.00$/0.000.00 / 0.00 0.00/$00/0.0$ 0.00
Rand.Per. (150) 0.00/$00/0.0$ 0.00 $\underline{0.65/0.22}$ 0.00/$00/0.0$ 0.00 $0.02/0.00$ 0.00/$00/0.0$ 0.00
Synthetic Data
We perform a numerical analysis based on the simple two-squares dataset [1] (drift intensity 5). We create the adversarial streams using Algorithm 2 where line 6 is performed by hand to assure $v_{i}\in\{0,1\}$ with as little changes as possible (see Table 1). Each stream has a length of $1{,}000$ samples, window sizes of the sliding window is 100, and (initial) reference window is 100/150. We use the permutation MMD test [1] with $2{,}500$ permutations and consider the smallest $p$-value found in the stream. We performed $500$ independent runs for each setup. The results are reported in Table 2. As can be seen, there is a (nearly) perfect alignment of our theoretical predictions and the empirical results with only one exceptional case.
Application to Water Distribution Networks
[…]
Refer to captionRefer to caption
[…]
5 Conclusion
In this paper, we considered the concept of drift adversarials. We showed that many commonly used drift detectors are subject to at least some drift adversarial attacks. We considered the problem from a general theoretical and concrete point of view and evaluated our findings empirically. Furthermore, we investigated the potential of our theory to construct problem-tailored drift detectors which seems to be a promising approach but requires further research.
Our considerations show that drift adversarials pose a major problem but might be numerically unstable. A further analysis is yet subject to future work.
Appendix A Proofs
In the following section, we will provide formal proofs of the statements made above.
A.1 Proof of Theorem 1
We start with the proof our the main theorem:
Proof of Theorem 1.
We have $d(\mathcal{D}_{W_{1}},\mathcal{D}_{W_{2}})=0$ if and only if $\mathcal{D}_{W_{1}}=\mathcal{D}_{W_{2}}$ if and only if
00 0 $=\|\mathcal{D}_{W_{1}}-\mathcal{D}_{W_{2}}\|_{\textnormal{TV}}$
$=\sup_{S\in\Sigma_{\mathcal{X}}}\left|\mathcal{D}_{W_{1}}(S)-\mathcal{D}_{W_{2}}(S)\right|$
$=\sup_{S\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{1})^{-1}\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{2})^{-1}\int\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|$
$=\frac{\sup_{S\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|}{P_{T}(W_{1})P_{T}(W_{2})}$
and therefore
00 0 $=\sup_{(W_{1},W_{2})\in\mathcal{W}}d(\mathcal{D}_{W_{1}},\mathcal{D}_{W_{2}})$
$\Leftrightarrow 0$ $=\sup_{(W_{1},W_{2})\in\mathcal{W}}\|\mathcal{D}_{W_{1}}-\mathcal{D}_{W_{2}}\|_{\textnormal{TV}}$
$\Leftrightarrow 0$ $=\sup_{(W_{1},W_{2})\in\mathcal{W}}\sup_{S\in\Sigma_{\mathcal{X}}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|$
$=\sup_{S\in\Sigma_{\mathcal{X}}}\sup_{(W_{1},W_{2})\in\mathcal{W}}\left|P_{T}(W_{2})\int_{W_{1}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)-P_{T}(W_{1})\int_{W_{2}}\mathcal{D}_{t}(S)\textnormal{d}P_{T}(t)\right|$
which is the case if and only if $t\mapsto\mathcal{D}_{t}(S)\in\textnormal{Adv}_{0}(A)$.
Obviously $\textnormal{Adv}_{0}(A)$ is convex. So in particular, if $f\in\textnormal{Adv}_{0}(A)$ then $\lambda f+\mu\in\textnormal{Adv}_{0}(A)$ if the resulting function stays in $[0,1]$. Hence for any $P,Q\in\textnormal{Pr}(\mathcal{X})$ and $S\in\Sigma_{\mathcal{X}}$ we have $f(t)P(S)+(1-f(t))Q(S)=f(t)(P(S)-Q(S))+Q(S)\in\textnormal{Adv}_{0}(A)$.
Conversely, $\mathcal{D}_{t}$ has drift if and only if $t\mapsto\mathcal{D}_{t}(S)\not\in\text{Const}(\mathcal{T})$ for some $S\in\Sigma_{\mathcal{X}}$. As $A$ detects the drift if and only if $t\mapsto\mathcal{D}_{t}(S)\not\in\textnormal{Adv}_{0}(A)$ for some $S\in\Sigma_{\mathcal{X}}$ and $\textnormal{Adv}_{0}(A)=\text{Const}(\mathcal{T})\cup\textnormal{Adv}(A)$ we see that the statement “$A$ detects drift if and only if $\mathcal{D}_{t}$ has drift” holds true if and only if $\textnormal{Adv}(A)=\emptyset$. ∎
A.2 Proofs of the statements in Table 1
We will now compute the improper adversarial sets displayed in Table 1. To do so, we will compute the sets
$\ker\mathcal{W}=\left\{f\in L^{1}_{\textnormal{loc}}(\mathcal{T})\>\left|\>\lambda(W_{2})\int_{W_{1}}f(t)\textnormal{d}t=\lambda(W_{1})\int_{W_{2}}f(t)\textnormal{d}t\forall(W_{1},W_{2})\in\mathcal{W}\right.\right\}.$
Proposition 1 (Fixed reference).
For $\mathcal{W}=\{([0,a],[t,t+l])\mid a\leq t\}$ with $a,l>0$ and $f\in L^{1}_{\textnormal{loc}}(\mathbb{R}_{\geq 0})$ the following are equivalent
  1. 1.
    $f\in\ker\mathcal{W}$
  2. 2.
    for all $t>a$ we have
    $\frac{l}{a}\int_{0}^{a}f(x)\textnormal{d}x=\int_{t}^{t+l}f(x)\textnormal{d}x$
  3. 3.
    $f$ is $l$-periodic after $a$, i.e., $f(t)=f(t+l)\forall t>a$, and has the same mean as before, i.e., $\frac{l}{a}\int_{0}^{a}f(x)\textnormal{d}x=\int_{a}^{a+l}f(x)\textnormal{d}x$.
In particular, $\textnormal{Adv}_{0}(\mathcal{W})\mathbb{R}=\ker\mathcal{W}$ are the functions that are $l$-periodic after $a$ and have the same mean as before. There are no boundary effects.
Proof.
1. $\Leftrightarrow$ 2. is easily seen by rewriting.
2. $\Rightarrow$ 3. define $C\in\mathbb{R}$ and $F:\mathbb{R}_{\geq 0}\times\mathbb{N}\to\mathbb{R}$ as
$C$ $:=\frac{l}{a}\int_{0}^{a}f(x)\textnormal{d}x$ and
$F_{t}(n)$ $:=\int_{0}^{a+t+ln}f(x)\textnormal{d}x.$
By definition of $C,F_{t}$ and by assumption we have $C=F_{t}(n+1)-F_{t}(n)$ for all $t\geq 0$ and $n\in\mathbb{N}$. Considering this as a recurrent equation in $n$ for every single $t$ it thus follows
$F_{t}(n)=c(t)+(t/l+n)\cdot C$
where $c(t)$ is the function parameter for each $t$ which is uniquely determined by the equation.
On the other hand, since $F_{t+l}(n)=F_{t}(n+1)$ and $(t+l)/l+n=t/l+(n+1)$ we have $c(t)=c(t+l)$ so $c$ is $l$-periodic. Therefore, the difference
$F_{t+h}(0)-F_{t}(0)$ $=(c(t+h)-c(t))+h/lC$
$=(c(t+h+l)-c(t+l))+h/lC$
$=F_{t+h+l}(0)-F_{t+l}(0)$
is $l$-periodic, too. On the other hand by Lebesgue’s differentiation theorem $(F_{t+h}(0)-F_{t-h}(0))/(2h)\to f(t)$ as $h\to 0$ for almost every $t$. Therefore, $f$ is $l$-periodic as well for all $t\geq a$. Furthermore, we can conclude that
$\int_{a}^{a+l}f(x)\textnormal{d}x=\frac{l}{a}\int_{0}^{a}f(x)\textnormal{d}x.$
3. $\Rightarrow$ 2. if $f$ is $l$-periodic after $a$ we have $\int_{t}^{t+l}f(x)\textnormal{d}x=c$ is $t$-invariant for all $t>a$ and if it has the same mean it holds
$\frac{l}{a}\int_{0}^{a}f(x)\textnormal{d}x=\int_{a}^{a+l}f(x)\textnormal{d}x=\int_{t}^{t+l}f(x)\textnormal{d}x.$
Therefore, the statement follows. ∎
To prove the next statement we need the following simple lemma:
Lemma 1.
Let $a_{n},b_{n}\subset\mathbb{R}$ and $c_{1},c_{2}\in\mathbb{R}$ with $c_{1}\neq c_{2}$. The following are equivalent
  1. 1.
    The limits $a_{n}$ and $b_{n}$ as $n\to\infty$ exist
  2. 2.
    The limits $a_{n}+c_{1}b_{n}$ and $a_{n}+c_{2}b_{n}$ exist as $n\to\infty$
Proof.
1. $\Rightarrow$ 2. is just linearity.
2. $\Rightarrow$ 1. follows by considering $(a_{n}-c_{1}b_{n})-(a_{n}-c_{2}b_{n})=(c_{1}-c_{2})b_{n}$. As the limit $n\to\infty$ on the left hand side exists so does the limit on the right hand side which by linearity implies that the limit of $b_{n}$ exists. Thus, by linearity, the limit of $a_{n}$ exists, too. ∎
Proposition 2 (Sliding windows).
For $\mathcal{W}=\{([t-l,t],[t,t+l])\mid t\in\mathbb{R}\}$ with $l>0$ and $f\in L^{1}_{\textnormal{loc}}(\mathbb{R})$ the following are equivalent
  1. 1.
    $f\in\ker\mathcal{W}$
  2. 2.
    for all $t\in\mathbb{R}$ we have
    $\int_{t-l}^{t}f(x)\textnormal{d}x=\int_{t}^{t+l}f(x)\textnormal{d}x$
  3. 3.
    $f(x)=p(x)+tq(x)$ with $p$ and $q$ $l$-periodic and $\int_{0}^{l}q(x)\textnormal{d}x=0$.
In particular, the solution has boundary effects, and $\textnormal{Adv}_{0}(\mathcal{W})\mathbb{R}\subsetneq\ker\mathcal{W}$ are the $l$-periodic functions.
Proof.
1. $\Leftrightarrow$ 2. is easily seen by rewriting.
2. $\Rightarrow$ $f(t)=q(t)+tq(t)$ with $p,q$ $l$-periodic: defining $F_{t}(n)=\int_{0}^{t+nl}f(x)\textnormal{d}x$ induces the a recurrent equation for every $t$
$F_{t}(n-1)+F_{t}(n+1)-2F_{t}(n)=0$
which has the solution $F_{t}(n)=c_{1}(t)+c_{2}(t)(t/l+n)$ with parameter functions $c_{1},c_{2}$.
As $F_{t+ml}(n)=F_{t}(n+m)$ we have
00 0 $=F_{t+ml}(n)-F_{t}(n+m)$
$=(c_{1}(t+ml)-c_{1}(t))+(c_{2}(t+ml)-c_{2}(t))(t/l+n+m)$
and therefore
00 0 $=0-0$
$=(F_{t+ml}(n)-F_{t}(n+m))-(F_{t+ml}(n+1)-F_{t}(n+m+1))$
$=(c_{2}(t+ml)+c_{2}(t))$
and hence $c_{2}$ is $l$-periodic but this then also implies that
$0=(c_{1}(t+ml)-c_{1}(t))+\underbrace{(c_{2}(t+ml)-c_{2}(t))}_{=0}(t/l+n+m)$
so $c_{1}$ is $l$-periodic, too.
Furthermore, as $t\mapsto F_{t}(n)$ is continuous and we have
$|c_{2}(t+h)-c_{2}(t)|$ $=|(F_{t+h}(0)-F_{t}(0))-(F_{t+h}(1)-F_{t}(1))|$
$\leq|F_{t+h}(0)-F_{t}(0)|+|F_{t+h}(1)-F_{t}(1)|\xrightarrow{h\to 0}0$
we conclude that $c_{2}$ and hence $c_{1}$ are continues, too.
Now considering
$\frac{F_{t+h}(0)-F_{t-h}(0)}{2h}$ $=\quad\frac{c_{1}(t+h)-c_{1}(t-h)}{2h}$
$\quad+\frac{c_{2}(t+h)-c_{2}(t-h)}{2h}\cdot(t+h)$
$\quad+c_{2}(t+h)\cdot\frac{(t+h)-(t-h)}{2h}$
$=\quad\underbrace{\frac{c_{1}(t+h)-c_{1}(t-h)}{2h}+c_{2}(t+h)}_{=:P(t)}$
$\quad+t\cdot\underbrace{\frac{c_{2}(t+h)-c_{2}(t-h)}{2h}}_{=:Q(t)}$
$\quad+\underbrace{\frac{c_{2}(t+h)-c_{2}(t-h)}{2}}_{\in o(h)}$
we see the quotient is of the form $P(t)+tQ(t)+o(h)$ for all $h>0$ with $P$ and $Q$ $l$-periodic since $c_{1}$ and $c_{2}$ are $l$-periodic. On the one hand, have $(F_{t+h}(0)-F_{t-h}(0))/(2h)\to f(t)$ as $h\to 0$ by Lebesgue’s differentiation theorem. As $P$ and $Q$ are $l$-periodic we obtain the same sequences for $a_{n}:=P(t+h_{n})=P(t+h_{n}+l)$ and $b_{n}:=Q(t+h_{n})=Q(t+h_{n}+l)$ so by Lemma 1 with $c_{1}=t$ and $c_{2}=t+l$ we see that the (point-wise) limits exist and thus obtain $l$-periodic $L^{1}_{\textnormal{loc}}$-function $p$ and $q$ with $f(t)=p(t)+q(t)t$.
As $f(t)\in[0,1]$ it follows that $q(t)=0$ as otherwise $q(t)t>1$ for $|t|$ sufficiently large. In particular, the solution has boundary effects.
2. $\Leftrightarrow$ 3. if 3. holds or – according to the previous claim – if 2. holds, we have $f(t)=p(t)+tq(t)$ with $p,q\in L^{1}_{\textnormal{loc}}(\mathbb{R})$ $l$-periodic. Therefore, it remains to show that $3.\Leftrightarrow\int_{0}^{l}q(x)\textnormal{d}x=0$. It holds
$\int_{t-l}^{t}f(x)\textnormal{d}x-\int_{t}^{t+l}f(x)\textnormal{d}x$ $=\int_{t-l}^{t}p(x)+xq(x)\textnormal{d}x-\int_{t}^{t+1}p(x)+xq(x)\textnormal{d}x$
$=\int_{t-l}^{t}p(x)+xq(x)\textnormal{d}x-\int_{t-l}^{t}\underbrace{p(x-l)}_{=p(x)}+(x-l)\underbrace{q(x-l)}_{=q(x)}\textnormal{d}x$
$=\int_{t-l}^{t}((x)-(x-l))q(x)\textnormal{d}x$
$=l\int_{t-l}^{t}q(x)\textnormal{d}x$
and as $q$ is $l$-periodic we have $\int_{t-l}^{t}q(x)\textnormal{d}x=\int_{0}^{l}q(x)\textnormal{d}x$. Therefore $f\in\ker\mathcal{W}$ if and only if $\int_{0}^{l}q(x)\textnormal{d}x=0$. ∎
Proposition 3 (Growing reference).
Let $\mathcal{W}=\{([0,t],[t,t+l])\mid t\geq a\}$ with $a,l>0$ and $f\in L^{1}_{\textnormal{loc}}(\mathbb{R}_{\geq 0})$. The following are equivalent:
  1. 1.
    $f\in\ker\mathcal{W}$
  2. 2.
    for all $t>a$ we have
    $\frac{l}{t}\int_{0}^{t}f(x)\textnormal{d}x=\int_{t}^{t+l}f(x)\textnormal{d}x$
  3. 3.
    $f(t)=\mathbb{I}_{[0,a]}(t)g(t)+p(t)+tq(t)$ for all $t>0$ with $p$ and $q$ $l$-periodic, $p(t)=\int_{0}^{t}q(x)\textnormal{d}x+C$, $\int_{0}^{l}q(x)\textnormal{d}x=0$, and $\int_{0}^{a}g(x)\textnormal{d}x=0$.
In particular, there are boundary effects, i.e., $\textnormal{Adv}_{0}(\mathcal{W})\mathbb{R}\subsetneq\ker\mathcal{W}$, and $\textnormal{Adv}_{0}(\mathcal{W})\mathbb{R}$ are exactly the functions that are arbitrary before $a$ and then constant with the same mean as before, i.e., $a^{-1}\int_{0}^{a}f(x)\textnormal{d}x=f(t)$ for all $t>a$.
Proof.
1. $\Leftrightarrow$ 2. is easily seen by rewriting.
2. $\Rightarrow$ Properties of 3. except $\int_{0}^{a}g(x)\textnormal{d}x=0$ define $F_{t}(n)=\int_{0}^{t+ln}f(x)\textnormal{d}x$ then we have
$(t+l(n-1))^{-1}F_{t}(n-1)=l^{-1}(F_{t}(n)-F_{t}(n-1))$
for all $t\geq a$ which allows the recursive definition for each $t>a$
$F_{t}(n)$ $=\frac{l+l(n-1)+t}{l(n-1)+t}F_{t}(n-1)$
$=\underbrace{\frac{ln+t}{l(n-1)+t}\frac{l(n-1)+t}{l(n-2)+t}}_{=\frac{ln+t}{l(n-2)+t}}F_{t}(n-2)$
$=\cdots=\frac{ln+t}{t}F_{t}(0)$
$=(t+ln)\underbrace{\frac{1}{t}F_{t}(0)}_{=:c(t)}$
As before we conclude from $F_{t+l}(n)=F_{t}(n+1)$ that
$0=F_{t}(n+1)-F_{t+l}(n)=(t+l(n+1))(c(t)-c(t+l))$
so $c$ is $l$-periodic. And from
$(F_{t+h}(1)-F_{t}(1))-(F_{t+h}(0)-F_{t}(0))$ $=c(t+h)-c(t)$
that since $t\mapsto F_{t}(n)$ is continuous that $c(t)$ is continuous. We extend $c$ to a function on all of $\mathbb{R}$ by using its periodicity.
[…]
$\frac{F_{t+h}(0)-F_{t-h}(0)}{2h}$ $=\left(t-h\right)\frac{c(t+h)-c(t-h)}{2h}+c(t+h).$
Thus, by Lebesgue’s differentiation theorem, we then conclude that $f(t)=\mathbb{I}_{[0,a]}(t)g(t)+p(t)+tq(t)$ with $p,q$ $l$-periodic, and $p(t)=C+\int_{0}^{t}q(x)\textnormal{d}x$ for all $t>0$ and $g$ arbitrary. In particular, as $p$ is $l$-periodic we have $0=p(l)-p(0)=\int_{0}^{l}q(x)\textnormal{d}x$.
2. $\Leftrightarrow$ 3. due to the relations, we have
$Q(t)$ $=\int_{0}^{t}q(x)\textnormal{d}x=p(t)-C,$ and
$\int_{0}^{t}Q(x)\textnormal{d}x$ $=\int_{0}^{t}p(x)\textnormal{d}x-tC.$
Denote by $C_{0}=\int_{0}^{a}g(x)\textnormal{d}x$, then we have
$\frac{1}{t}\int_{0}^{t}f(x)\textnormal{d}x$ $=\frac{1}{t}\left(\int_{0}^{a}g(x)\textnormal{d}x+\int_{0}^{t}p(x)+xq(x)\textnormal{d}x\right)$
$=\frac{1}{t}\left(C_{0}+\int_{0}^{t}p(x)\textnormal{d}x+\left[Q(x)x\right]_{0}^{t}-\int_{0}^{t}Q(x)\textnormal{d}x\right)$
$=\frac{1}{t}\left(C_{0}+tC+Q(t)t\right)$
$=\frac{C_{0}}{t}+C+Q(t)$
$\frac{1}{l}\int_{t}^{t+l}f(x)\textnormal{d}x$ $=\frac{1}{l}\left(\int_{t}^{t+l}Q(x)\textnormal{d}x+lC+\left[Q(x)x\right]_{t}^{t+l}-\int_{t}^{t+l}Q(x)\textnormal{d}x\right)$
$=\frac{1}{l}\left(lC+Q(t)l\right)$
$=C+Q(t),$
$\Rightarrow\frac{C_{0}}{t}$ $=\frac{1}{t}\int_{0}^{t}f(x)\textnormal{d}x-\frac{1}{l}\int_{t}^{t+l}f(x)\textnormal{d}x$
we conclude that $f\in\ker\mathcal{W}$ if and only if $C_{0}=0$.
As before we conclude that form $|f(t)|\leq 1$ that $q(t)=0$ and hence $p(t)=C$ is constant and $a^{-1}\int_{0}^{a}g(x)+p(x)\textnormal{d}x=l^{-1}\int_{a}^{a+l}p(x)\textnormal{d}x=C$ with $-C\leq g(t)\leq 1-C$. ∎